# A Novel Data Augmentation Approach for Automatic Speaking Assessment on Opinion Expressions

### v1

@@ 1 Introduction
More recently, multi-task and multimodal architectures have been introduced to capture complementary information inherent in both the text content and speech signal of a spoken response. The emergence of self-supervised speech encoders (e.g., wav2vec 2.0 [5, 7, 17]), Transformer-based text models [18], and multimodal large language model (MLLM) such as Phi-4 multimodal  [^1] Microsoft Phi-4 multimodal https://arxiv.org/abs/2503.01743 gives rise to a variety of end-to-end modeling frameworks that learn response representations jointly from raw audio and textual transcripts, markedly improving scoring accuracy and generalization. Nevertheless, these models still rely excessively on large amounts of annotated training datasets that are frequently scarce for ASA on opinion expressions, thus often leading to overfitting on limited prompts.
@@
To alleviate data scarcity, prior studies explore combining text-to-speech (TTS) techniques with LLM–driven data augmentation. These approaches have shown some success on various speech and language processing tasks, such as automatic speech recognition (ASR) [19, 20], emotion recognition [21], text classification [22, 23], and intent recognition [24]. For the purpose of spoken content generation, these methods typically use LLMs to generate diverse texts, thereby increasing content variety, and subsequently employ TTS systems to synthesize speech from the generated texts, enriching speaker characteristics, prosody, and content. Because synthesized and real texts and audio waves differ in their acoustic feature distributions, there is good reason to apply dynamic instance reweighting or metric-learning techniques to adaptively adjust the importance of each synthesized instance based on its proximity to real counterparts in a embedding space, thus narrowing the domain gap [22]. However, to our knowledge, there is still a dearth of work investigating the synergistic effect of combining the aforementioned methods and their extensions to automated scoring of spoken opinion expressions.
@@
In view of this, we present a unified framework for spoken opinion‐expression assessment under the constraint of limited resources, which combines generative data augmentation with neural multimodal modeling. First, we develop a two‐step augmentation pipeline in which an LLM generates proficiency‐conditioned text responses [25] and a high‐fidelity TTS system synthesizes these text responses into spoken utterances while retaining non‐native prosody and natural disfluencies [20]. Next, we introduce a novel dynamic importance loss that builds on a recently proposed data‐weighting principle [22] for training of our scoring model. This training loss automatically increases the contribution of synthesized instances, so as to bridge domain gaps or proficiency distribution imbalance and meanwhile curb overreliance on the generated data. Finally, the backbone of our scoring model is a Phi‐4 multimodal 11footnotemark: 1 finetuned with LoRA adapters [26], which directly fuses text prompt, as well as raw audio and ASR transcript of a spoken response, to predict the corresponding proficiency score. Our approach delivers robust performance on both seen and unseen prompts despite that only limited labeled instances are made available. Our approach not only attains superior performance on low-resource speaking assessment but also offers a replicable paradigm for other multimodal, low-resource speaking assessment tasks. In summary, our key contributions are at least three-fold:
@@ 2 Methodology
Our methodology advances low-resource opinion-expression assessment through three key innovations. First, we introduce a proficiency-aware text synthesis module that leverages an LLM to generate response texts closely aligned with each proficiency level. This module constructs in-context prompts by sampling and permuting transcripts of real test-takers’ responses, ensuring that generated responses faithfully mimic authentic test-takers’ language competence. Second, we design a speaker-conditioned voice-cloning pipeline, leveraging embeddings extracted from real recordings to guide a multi-speaker TTS model to produce synthesized utterances that preserve non-native prosody and natural hesitations. Third, we design and implement a novel dynamic importance loss, which automatically adjusts the importance weight of each training instance based on the relative confidence of a fixed “quality” model and the trainable “target” model. This concept parallels the data-weighting idea of Kuo et al. in text data generation [22], but our formulation is the first to integrate it into a fully end-to-end multimodal scoring network. We implement this loss within our scoring model, a Phi-4 multimodal 11footnotemark: 1, enhanced with LoRA adapters [26] for parameter-efficient fine-tuning. Together, these three innovations—rich text augmentation, authentic speech cloning, and principled loss reweighting—curate a training dataset that balances the diversity and fidelity factors between synthesized and real data, yielding robust scoring performance under stringent data constraints.
@@ 2.2 Voice Cloning Synthesis
When converting a synthesized texts $\hat{R_{l,i}}$ into the corresponding realistic speech of a test-taker, we observed that a standard TTS model tends to produce over-perfect audio which lacks natural characteristics of of non-native speakers, such as disfluencies, mispronunciations, and prosodic variability, since it is typically trained on corpora of native speakers. To remedy this, we employ a multi-speaker TTS conditioned on disparate speaker embeddings rather than a single fixed voice.
@@ 2.3 End-to-End Multimodal Training with Dynamic Importance Reweighting
After preparing synthesized texts and cloned speech, we introduce a new dynamic importance reweighting loss that automatically adjusts the weigh of each training instance according to its fidelity to real learner behavior. This loss computes the ratio of confidence scores produced by a fixed “quality model” and by our trainable “target model,” then uses that ratio to amplify high‐quality synthesized instances and suppress poorly matched ones. At the same time, real instances receive an elevated base weight to ensure the model remains grounded in authentic speech distributions. We integrate this loss into a neural scoring model built upon the Phi-4 multimodal 11footnotemark: 1 backbone and fine-tuned with LoRA adapters [26] for maximal parameter efficiency. The training phase unfolds in two stages:
Stage 1. Quality Model Calibration We first fine-tune the Phi-4 multimodal 11footnotemark: 1 backbone on real prompt–audio–score triplets alone, using standard cross-entropy to obtain a Quality Model $f_{Q}$. This model establishes a reliable reference distribution over genuine responses of test-takers.
@@
which will be employed to amplify those synthesized instances that most closely match real-data behavior. To prevent overreliance on generated data—whose fluency often exceeds natural learner speech—we introduce a real-instance weight $\alpha>1$. The combined per-instance loss is
$\ell_{i}=-\,r_{i}\,w_{i}\,\log\bigl{(}p_{T,i}\bigr{)},\quad w_{i}=\begin{cases}\alpha,&x_{i}\in\mathcal{D}_{\text{real}},\\ 1,&x_{i}\in\mathcal{D}_{\text{syn}}.\end{cases}$ (7)
@@
Acc. Binary Acc. Acc. Binary Acc.
@@ 3.1 Dataset
We used the LTTC GEPT Intermediate–Advanced opinion‐expression dataset, which comprises 1,200 prompt–response–score triples evenly drawn from four prompt types. Among them, 300 triples of the fourth type served as an unseen test set; the remaining 900 triples were split 8:1:1 into training, validation and seen-test sets. Each spoken response was rated 1–5 by two experts; we average and floor their scores (no responses scored 1), treating 4–5 as passing and 1–3 as failing. All recordings of spoken responses were manually transcribed. We then generated a synthesized corpus matching the size and score distribution of the training set, with utterance lengths varying according to the generated text. Figure 3 illustrates the final score distributions.
@@ 3.2 Experimental Setup
We generate synthesized texts with OpenAI o4-mini  [^2] OpenAI o4-mini https://platform.openai.com/docs/models/o4-mini using up to ten shuffled in-context examples per prompt and decoding at temperature $\tau=1.5$. For each text, we sample a real reference clip at the same proficiency level, extract its 512-D speaker embedding via Coqui-ai XTTSv2  [^3] Coqui-ai XTTSv2 https://github.com/coqui-ai/TTS, and synthesize learner-like speech with a multi-speaker TTS. Our scoring backbone is a Phi-4 multimodal 11footnotemark: 1 equipped with LoRA adapters [26]. We first fine-tune a Quality Model on real prompt–speech–score triplets for three epochs using AdamW ($\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-7}$), gradient clipping at $1.0$, learning rate $\alpha=2$, batch size $16$. We then train the Target Model on the combined real and synthesized corpus under the same settings, applying a real-sample weight $\alpha=2$ in our dynamic importance loss.
@@ 3.3 Experiment Results
Table 1 compares our proposed model with strong single-modality and dual-modality baselines on both the seen and unseen prompts. The baselines include an acoustics-only wav2vec2 [17] system, a text only BERT [18] model, and a Phi‐4 multimodal 11footnotemark: 1 finetuned on the real data alone.
On the test set of seen prompts, our model achieves the highest overall accuracy and pass–fail accuracy among all methods. By combining speech and text through LLM-driven augmentation and dynamic reweighting, it alleviate the downsides the systems that rely either on purely acoustic or purely textual information. It notably outperforms wav2vec in grading consistency and robustness, while surpasseing BERT despite BERT’s strong contextual encoding of transcripts. Although fine-tuning Phi‐4 multimodal 11footnotemark: 1 model on real data alone yields competitive results, our pipeline with synthesized augmentation and the dynamic importance reweighting loss delivers the best performance, confirming the practical effectiveness of our contributions.
On the the test et of unseen prompt, our model maintains its advantage and incurs a smaller performance drop than any baseline. The synergy of diverse synthesized data and adaptive loss weighting enables more dependable scoring of novel prompts, alleviating under- or over-scoring on unfamiliar content. These results show that our unified training framework not only boosts accuracy on familiar prompts but also enhances robustness across different prompts.

### v2

@@ 1 Introduction
More recently, multi-task and multimodal architectures have been introduced to capture complementary information inherent in both the text content and speech signal of a spoken response. The emergence of self-supervised speech encoders (e.g., wav2vec 2.0 [5, 7, 17]), Transformer-based text models [18], and multimodal large language model (MLLM) such as Phi-4 multimodal [19] gives rise to a variety of end-to-end modeling frameworks that learn response representations jointly from raw audio and textual transcripts, markedly improving scoring accuracy and generalization. Nevertheless, these models still rely excessively on large amounts of annotated training datasets that are frequently scarce for ASA on opinion expressions, thus often leading to overfitting on limited prompts.
@@
To alleviate data scarcity, prior studies explore combining text-to-speech (TTS) techniques with LLM–driven data augmentation [20]. These approaches have shown some success on various speech and language processing tasks, such as automatic speech recognition (ASR) [21, 22], emotion recognition [23], text classification [24, 25], and intent recognition [26]. For the purpose of spoken content generation, these methods typically use LLMs to generate diverse texts, thereby increasing content variety, and subsequently employ TTS systems to synthesize speech from the generated texts, enriching speaker characteristics, prosody, and content. Because synthesized and real texts and audio waves differ in their acoustic feature distributions, there is good reason to apply dynamic instance reweighting or metric-learning techniques to adaptively adjust the importance of each synthesized instance based on its proximity to real counterparts in a embedding space, thus narrowing the domain gap [24]. However, to our knowledge, there is still a dearth of work investigating the synergistic effect of combining the aforementioned methods and their extensions to automated scoring of spoken opinion expressions.
@@
In view of this, we present a unified framework for spoken opinion‐expression assessment under the constraint of limited resources, which combines generative data augmentation with neural multimodal modeling. First, we develop a two‐step augmentation pipeline in which an LLM generates proficiency‐conditioned text responses [27] and a high‐fidelity TTS system synthesizes these text responses into spoken utterances while retaining non‐native prosody and natural disfluencies [22]. Next, we introduce a novel dynamic importance loss that builds on a recently proposed data‐weighting principle [24] for training of our scoring model. This training loss automatically increases the contribution of synthesized instances, so as to bridge domain gaps or proficiency distribution imbalance and meanwhile curb overreliance on the generated data. Finally, the backbone of our scoring model is a Phi‐4 multimodal [19] finetuned with LoRA adapters [28], which directly fuses text prompt, as well as raw audio and ASR transcript of a spoken response, to predict the corresponding proficiency score. Our approach delivers robust performance on both seen and unseen prompts despite that only limited labeled instances are made available. Our approach not only attains superior performance on low-resource speaking assessment but also offers a replicable paradigm for other multimodal, low-resource speaking assessment tasks. In summary, our key contributions are at least three-fold:
@@ 2 Methodology
Our approach enhances opinion expression assessment in low resource settings through three tightly integrated innovations. First, a proficiency aware text synthesis module uses a large language model, prompted with randomly sampled and permuted transcripts of authentic test taker responses, to generate answers aligned with each proficiency level and faithful to real language use. Second, speaker conditioned voice cloning, driven by speaker embeddings extracted from real recordings, guides a multi speaker TTS model to convert these texts into speech while preserving non native prosody, pauses, and disfluencies; Figure 1 illustrates the complete speech synthesis pipeline. Finally, the proposed dynamic importance reweighting loss, shown in Figure 2, compares the confidences that a frozen quality model and a trainable target model assign to the correct label and then dynamically increases the weight of highly realistic synthetic samples while keeping genuine data predominant. By combining rich text augmentation, authentic speech cloning, and principled loss reweighting, our method yields a training corpus that balances diversity and authenticity and achieves robust scoring performance under strict data constraints. Implementation details and experimental results appear in the following sections.
@@ 2.2 Voice Cloning Synthesis
When converting a synthesized texts $\hat{R_{l,i}}$ into the corresponding realistic speech of a test-taker, we observed that a standard TTS model tends to produce over-perfect audio which lacks natural characteristics of non-native speakers, such as disfluencies, mispronunciations, and prosodic variability, since it is typically trained on corpora of native speakers. To remedy this, we employ a multi-speaker TTS conditioned on disparate speaker embeddings rather than a single fixed voice.
@@ 2.3 End-to-End Multimodal Training with Dynamic Importance Reweighting
After preparing synthesized texts and cloned speech, we introduce a new dynamic importance reweighting loss that automatically adjusts the weigh of each training instance according to its fidelity to real learner behavior. This loss computes the ratio of confidence scores produced by a fixed “quality model” and by our trainable “target model,” then uses that ratio to amplify high‐quality synthesized instances and suppress poorly matched ones. At the same time, real instances receive an elevated base weight to ensure the model remains grounded in authentic speech distributions. We integrate this loss into a neural scoring model built upon the Phi-4 multimodal [19] backbone and fine-tuned with LoRA adapters [28] for maximal parameter efficiency. The training phase unfolds in two stages:
Stage 1. Quality Model Calibration We first fine-tune the Phi-4 multimodal [19] backbone on real prompt–audio–score triplets alone, using standard cross-entropy to obtain a Quality Model $f_{Q}$. This model establishes a reliable reference distribution over genuine responses of test-takers.
@@
which will be employed to amplify those synthesized instances that most closely match real-data behavior. To prevent overreliance on generated data—whose fluency often exceeds natural learner speech—we introduce a real-instance weight $\alpha>1$ with learning rate $\eta$. The combined per-instance loss is
$\ell_{i}=-\,\eta\,r_{i}\,w_{i}\,\log\bigl{(}p_{T,i}\bigr{)},\quad w_{i}=\begin{cases}\alpha,&x_{i}\in\mathcal{D}_{\text{real}},\\ 1,&x_{i}\in\mathcal{D}_{\text{syn}}.\end{cases}$ (7)
@@
Overall Acc. Pass–Fail Acc. Overall Acc. Pass–Fail Acc.
@@ 3.1 Dataset
We used the LTTC GEPT Intermediate to Advanced opinion expression dataset, which contains 1 200 prompt–response–score triples evenly distributed across four prompt types within the discussion subtask. Of these, 300 triples belonging to the fourth prompt type were set aside as an unseen test set in order to simulate domain transfer. The remaining 900 triples were divided in an 8:1:1 ratio into training, validation, and seen-test sets. Each spoken response received a 1-to-5 score from two expert raters. Following the official GEPT standard, we averaged the two scores, then applied floor rounding (no response ended with a score of 1). Scores of 4 or 5 were labeled passing and scores of 1 to 3 were labeled failing; this binary label reflects the GEPT Intermediate to Advanced speaking threshold. All spoken responses were manually transcribed for model input. To alleviate data scarcity, we generated a synthesized corpus that matches the size and score distribution of the real training set. Figure 3 shows the final score distributions for real data.
@@ 3.2 Experimental Setup
We generate synthesized texts with OpenAI o4-mini  [^1] OpenAI o4-mini https://platform.openai.com/docs/models/o4-mini using up to ten shuffled in-context examples per prompt and decoding at temperature $\tau=1.5$. For each text, we sample a real reference clip at the same proficiency level, extract its 512-D speaker embedding via Coqui-ai XTTSv2  [^2] Coqui-ai XTTSv2 https://github.com/coqui-ai/TTS, and synthesize learner-like speech with a multi-speaker TTS. Our scoring backbone is a Phi-4 multimodal [19] equipped with LoRA adapters [28]. We first fine-tune a Quality Model on real prompt–speech–score triplets for three epochs using AdamW ($\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-7}$), gradient clipping at $1.0$, learning rate $\alpha=2$, batch size $16$. We then train the Target Model on the combined real and synthesized corpus under the same settings, applying a real-sample weight $\alpha=2$ in our dynamic importance loss.
@@ 3.3 Experiment Results
Table 1 compares our proposed model with strong single-modality and dual-modality baselines on both the seen and unseen prompts. The baselines include an acoustics-only wav2vec2 [17] system, a text only BERT [18] model, and a Phi‐4 multimodal [19] finetuned on the real data alone.
On the test set of seen prompts, our model achieves the highest overall accuracy and pass–fail accuracy among all methods. By combining speech and text through LLM-driven augmentation and dynamic reweighting, it alleviate the downsides the systems that rely either on purely acoustic or purely textual information. It notably outperforms wav2vec in grading consistency and robustness, while surpasseing BERT despite BERT’s strong contextual encoding of transcripts. Although fine-tuning Phi‐4 multimodal model on real data alone yields competitive results, our pipeline with synthesized augmentation and the dynamic importance reweighting loss delivers the best performance, confirming the practical effectiveness of our contributions.
On the the test of unseen prompt, our model maintains its advantage and incurs a smaller performance drop than any baseline. The synergy of diverse synthesized data and adaptive loss weighting enables more dependable scoring of novel prompts, alleviating under- or over-scoring on unfamiliar content. These results show that our unified training framework not only boosts accuracy on familiar prompts but also enhances robustness across different prompts.